OpenStudy (cutiecomittee123):

How do I solve this equation for x? 6e^2x=7e^4x

OpenStudy (freckles):

\[7e^{4x}-6e^{2x}=0\] you can start by factoring

OpenStudy (freckles):

I make it easier to see \[7e^{2x+2x}-6e^{2x}=0 \\ 7e^{2x}e^{2x}-6e^{2x}=0\] you should see both the first and second term on the left hand expression have common factor

OpenStudy (cutiecomittee123):

I know the answer to the question is 0 but I just need to know how to get there.

OpenStudy (freckles):

do you know how to factor?

OpenStudy (cutiecomittee123):

Yeah I do

OpenStudy (freckles):

can you factor what I asked you to factor

OpenStudy (cutiecomittee123):

In this equation I am confused on how to factor though

OpenStudy (freckles):

do you see both terms have e^(2x)?

OpenStudy (cutiecomittee123):

Yeah I see that

OpenStudy (freckles):

\[e^{2x}(7e^{2x}-6)=0\] just factor it out then :p

OpenStudy (freckles):

set both factors equal to 0 and solve for x

OpenStudy (freckles):

\[e^{2x}=0 \text{ or } 7e^{2x}-6=0\] hint e^(2x) is never 0 first equation cannot be solved you just have to look at second equation now

OpenStudy (freckles):

\[e^{2x}=\frac{6}{7}\] can you see if you can solve from here?

OpenStudy (cutiecomittee123):

I cant. Maybe you could show me but explain it as you go

OpenStudy (freckles):

do you know that ln(e^x)=x?

OpenStudy (freckles):

just take ln( ) of both sides

OpenStudy (cutiecomittee123):

So ln(e^2x)/ln(6/7)

OpenStudy (cutiecomittee123):

Dont you just divide like so ^^^

OpenStudy (freckles):

you changed equal sign to division sign?

OpenStudy (cutiecomittee123):

Oh so it is just ln(e^2x) = ln(6/7)

OpenStudy (freckles):

\[e^{2x}=\frac{6}{7} \\ \text{ equal sign does \not turn into division sign \it is an equal sign } \\ \ln(e^{2x})=\ln(\frac{6}{7})\] good now use the fact that exp and ln are inverse functions of each other

OpenStudy (freckles):

that is I mean y=e^x and y=ln(x) are inverse of each other that is ln(e^x)=x or e^(ln(x))=x for x>0

OpenStudy (freckles):

or you can use power rule for log

OpenStudy (freckles):

that is ln(x^r)=rln(x)

OpenStudy (freckles):

you bring power of the inside down in front

OpenStudy (cutiecomittee123):

Woah wait so then I just go 2xln=ln6/7 ??? I am confused

OpenStudy (freckles):

well you are missing the inside of that log

OpenStudy (freckles):

what happen to the e

OpenStudy (freckles):

\[2x \ln(e)=\ln(\frac{6}{7})\]

OpenStudy (freckles):

by the way ln(e)=1 so you really just have 2x on the left hand side

OpenStudy (cutiecomittee123):

Oh gotcha

OpenStudy (cutiecomittee123):

So then what??

OpenStudy (freckles):

solve for x

OpenStudy (freckles):

you only have one last step

OpenStudy (cutiecomittee123):

divide by 2?

OpenStudy (freckles):

you know you would solve something like 2x=14 by dividing 2 on both sides yes exactly same thing here divide 2 on both sides

OpenStudy (cutiecomittee123):

Alright and I get x= -.077 as the answer :)

OpenStudy (freckles):

that is approximation

OpenStudy (freckles):

\[2x=\ln(\frac{6}{7}) \\ x=\frac{1}{2} \ln(\frac{6}{7})\]

OpenStudy (freckles):

that is exact if you wanted to know :p

OpenStudy (cutiecomittee123):

sweet can you help me with one more? This one will be a lot quicker.

OpenStudy (freckles):

ok i will try

OpenStudy (cutiecomittee123):

10=1.5e^x

OpenStudy (cutiecomittee123):

So I already attempted it and I just did ln(10)/ln(1.5)

OpenStudy (cutiecomittee123):

and I got approximately .934

OpenStudy (freckles):

one sec multitasking

OpenStudy (freckles):

ok

OpenStudy (freckles):

\[10=1.5e^x \\ \frac{10}{1.5}=e^x\]

OpenStudy (freckles):

as before just take ln( ) of both sides

OpenStudy (freckles):

and ln(a/b) doesn't equal ln(a)/ln(b)

OpenStudy (cutiecomittee123):

Okay so ln(10)=ln(1.5)

OpenStudy (freckles):

that isn't a true equation

OpenStudy (freckles):

e^x=a means x=ln(a) assuming of course a>0

OpenStudy (freckles):

your a is 10/1.5

OpenStudy (cutiecomittee123):

Okay so ln(10/1.5)

OpenStudy (freckles):

yep

OpenStudy (cutiecomittee123):

Then just put that into the calculator and solve??

OpenStudy (freckles):

\[a=b \cdot e^x \\ \text{ first step isolate } e^x \\ \text{ do this by dividing } b \text{ on both sides } \\ \frac{a}{b}=e^x \\ \text{ then take } \ln( ) \text{ of both sides } \\ \ln(\frac{a}{b})=\ln(e^x) \\ \ln(\frac{a}{b})=x \ln(e) \\ \text{ recall } \ln(e)=1 \\ \ln(\frac{a}{b})=x\] well I would leave it like this but yeah if you want an approximation

OpenStudy (cutiecomittee123):

So I calculated and got x is approximately 1.897

OpenStudy (freckles):

yep seems great

OpenStudy (cutiecomittee123):

sweet, thanks:)

OpenStudy (freckles):

np